OpenStudy (anonymous):

Solve x^2 - 9 = 0 x = 3 no real solution x = 3, -3 x = -3

OpenStudy (anonymous):

add9 to both sides, then take the square root of both sides. the answer is x=3

jimthompson5910 (jim_thompson5910):

keep in mind that if you have x^2 = 9, it would be x = 3 or x = -3

jimthompson5910 (jim_thompson5910):

another way to do it x^2 - 9 = 0 (x-3)(x+3) = 0 x-3=0 or x+3=0 x = 3 or x = -3